OpenStudy (anonymous):

The roof lines of a building can be described by the system of equations below, where the floor is represented by the x-axis and the y-axis and the height of the building is along the z-axis. Measurements are in feet. Find the point of intersection of the roof lines, using an ordered triple in the form (x, y, z) {x+y+z=53 3x-2+z=69 -x+2y-z=-59

OpenStudy (anonymous):

airricuhh shouldnt the second equation of a term with y

OpenStudy (anonymous):

**have a term with y**

OpenStudy (anonymous):

im sorry, its {x+y+z=53 3x-2y+z=69 -x+2y-z=-59

OpenStudy (anonymous):

x=5;y=-2and z=50

OpenStudy (anonymous):

can i ask you another question?